### v5

# A non-backtracking method for long matrix and tensor completion

Abstract
We consider the problem of low-rank rectangular matrix completion in the regime where the matrix $M$ of size $n\times m$ is “long”, i.e., the aspect ratio $m/n$ diverges to infinity. Such matrices are of particular interest in the study of tensor completion, where they arise from the unfolding of a low-rank tensor. In the case where the sampling probability is $\frac{d}{\sqrt{mn}}$, we propose a new spectral algorithm for recovering the singular values and left singular vectors of the original matrix $M$ based on a variant of the standard non-backtracking operator of a suitably defined bipartite weighted random graph, which we call a non-backtracking wedge operator. When $d$ is above a Kesten-Stigum-type sampling threshold, our algorithm recovers a correlated version of the singular value decomposition of $M$ with quantifiable error bounds. This is the first result in the regime of bounded $d$ for weak recovery and the first for weak consistency when $d\to\infty$ arbitrarily slowly without any polylog factors. As an application, for low-CP-rank orthogonal $k$-tensor completion, we efficiently achieve weak recovery with sample size $O(n^{k/2})$ and weak consistency with sample size $\omega(n^{k/2})$. A similar result is obtained for low-multilinear-rank tensor completion with $O(n^{k/2})$ many samples.
1 Introduction
Matrix completion is the problem of reconstructing a matrix from a (usually random) subset of entries, leveraging prior structural knowledge such as rank and incoherence. When the two dimensions $n,m$ of the matrix $M$ are comparable, this problem has been widely studied in the past decades, with a fairly general pattern emerging. Roughly speaking, if the ground truth matrix is low-rank and eigenvectors are sufficiently incoherent (or delocalized), sampling $\Omega(n\log n)$ many entries uniformly at random with a certain efficient optimization procedure, one can exactly recover the ground truth matrix [keshavan.montanari.ea_2009_matrix, keshavan2010matrix, candes.tao_2010_power, candes.plan_2010_matrix, recht2011simpler, candes2012exact, jain2013low]. This is not the case when the sample size is only $O(n)$, where full completion is provably impossible under uniform sampling. However, an approximation of the ground truth matrix is still possible [heiman2014deterministic, gamarnik2017matrix, brito2022spectral] with efficient algorithms, and more refined estimates of the eigenvectors were studied in [bordenave2020detection].
In recent years, attention has shifted to “long” matrix models where the row and column sizes of the matrix are not proportional. This is the case, for example, in the spiked long matrix and tensor models [montanari.wu_2022_fundamental, ding.hopkins.ea_2020_estimating, arous2021long, montanari.richard_2014_statistical]. This situation also occurs in bipartite graph clustering and community detection, where the goal is to recover the community structure in the smaller vertex set [florescu2016spectral, ndaoud2021improved, cai2021subspace, braun2022minimax, braun2023strong] for a bipartite stochastic block model, and in subspace recovery problems [cai2021subspace, zhou2021rate].
Our work is closely related to tensor completion, which is the higher-order analog of matrix completion. Indeed, tensor problems are highly related to their long matrix counterparts through the unfolding operation for an order-$k$ tensor of size $n^{k}$:
$\mathrm{unfold}_{a,b}:\mathbb{R}^{n^{k}}\to\mathbb{R}^{n^{a}\times n^{b}},$
where $a+b=k$. For example, an order-3 tensor of size $n$ yields an unfolded matrix of size $n\times n^{2}$; in general, uneven unfoldings (i.e., with $a\neq b$) of a tensor give rise to long rectangular matrices. In those tensor problems, the ground truth is assumed to have a low CP-rank, which implies that the unfolded matrix is low-rank. In low-rank $k$-tensor completion, the best known polynomial-time algorithms require $\tilde{O}(n^{k/2})$ revealed entries [jain2014provable, montanari.sun_2018_spectral, xia2019polynomial, cai2021nonconvex], while information theoretically $\tilde{O}(n)$ suffices [ghadermarzy2019near, harris2021, harris2023spectral]. In [barak.moitra_2016_noisy], the authors showed lower bounds for a family of algorithms based on the sum-of-squares hierarchy for order-3 noisy tensor completion, where $\Omega(n^{3/2})$ many samples are needed, by connecting it to the literature on refuting random 3-SAT. They also conjectured such a sample size is needed for any polynomial-time algorithm. This is aligned with other tensor estimation problems where a diverging statistical-to-computational gap is expected [montanari.richard_2014_statistical, wein2019kikuchi, jagannath2020statistical, arous2020algorithmic, dudeja2022statistical, auddy2021estimating]. However, so far no known work has reached the exact threshold $\Omega(n^{3/2})$, only approaching it up to polylog factors, in a similar fashion to the matrix completion phase diagram described above. For exact recovery results on tensor completion, all the methods in the literature are either convex optimization or spectral initialization with iterative optimization steps. Our method achieves weak consistency without any log factors, which could improve the sample size for spectral initialization in the work [jain2014provable, cai2021nonconvex, xia2019polynomial, xia2021statistically].
The applications we mentioned above are related to the same question:
Can we improve the sample complexity for estimating a one-sided structure in a long matrix?
Specifically, we explore the use of the non-backtracking operator in this context. The non-backtracking operator is an important object in the study of spectral graph theory [bass_iharaselberg_1992, terras2010zeta]. Recently, it has been a powerful tool for understanding the spectrum of a very sparse random matrix [bordenave2018nonbacktracking, stephan2020non, bordenave2020new, bordenave2019eigenvalues, brito2022spectral, benaych2020spectral, alt2021extremal, alt2021delocalization, dumitriu2022extreme, stephan2022sparse] and for designing efficient algorithms in community detection and matrix completion [krzakala2013spectral, bordenave2018nonbacktracking, stephan2020non, bordenave2020detection, stephan2022sparse, zhu2023non]. In the bipartite graph setting, the non-backtracking spectrum was analyzed in [brito2022spectral, dumitriu2022extreme, dumitriu2021spectra] when the two sides have a bounded aspect ratio. For the spectrum of the adjacency matrix of a random bipartite graph with a diverging aspect ratio, only a few results were obtained in [zhu2022second, dumitriu2020global] for random bipartite biregular graphs, and in [guruswami2022l_p] for the adjacency matrix of bipartite random graphs with random sign flips.
In this work, we aim to address this gap by defining a non-backtracking operator suitable for sparse matrix estimation with a diverging aspect ratio, which we call a non-backtracking wedge matrix. Our work sheds light on the fundamental limits and practical algorithms for sparse long matrix estimation with a focus on one-sided structure recovery. To the best of our knowledge, this is the first result for a long $n\times m$ low-rank matrix completion with sampling probability $p=\frac{d}{\sqrt{mn}}$ in the regime of bounded $d$ for weak recovery and the first result for weak consistency when $d\to\infty$ arbitrarily slowly without any polylog factors. These two types of recovery guarantees are well studied in the literature of community detection [abbe_2018_community]. In fact, a direct adaptation of the bipartite non-backtracking operator we define here leads to a simpler spectral algorithm for the detection of the bipartite stochastic block model studied in [florescu2016spectral]. Below, We summarize our results informally. See Theorems LABEL:thm:eigenvalues and LABEL:thm:eigenvectors for the precise statements.
Theorem 1 (long matrix completion, informal).
With $O(\sqrt{mn})$ many samples, if the left singular vectors of a long $n\times m$ matrix $(m\gg n)$ are sufficiently delocalized, and the singular values are above a specific Kesten-Stigum threshold, then the non-backtracking wedge operator can recover the ground truth singular values and find correlated estimates of the left singular vectors.
The Kesten-Stigum threshold in Theorem 1 is analogous to a similar threshold for community detection in the stochastic block model, which characterizes a threshold of the signal-to-noise ratio for the existence of polynomial time algorithms to achieve weak recovery; see [abbe_2018_community] for more details. We discuss the Kesten-Stigum bound in matrix completion in Section LABEL:sec:discussion.
We then apply our main results to the low-rank tensor completion problem under an orthogonality assumption on the $r$ components in a tensor decomposition. This is the first efficient algorithm to achieve weak recovery of a low-rank tensor with $\Omega(n^{k/2})$ samples without any extra $\mathrm{polylog}(n)$ factors. In particular, an explicit $n^{k/2}$ threshold is obtained for rank-1 tensor completion when the entries are ±$\pm 1$. The weak recovery results can serve as a good spectral initialization step for other optimization procedures [montanari.sun_2018_spectral, jain2014provable, cai2021nonconvex]. Unlike the common approach of unfolding a tensor to a matrix as square as possible ( i.e., applying $\mathrm{unfold}_{\lfloor k/2\rfloor,\lceil k/2\rceil}$ to $T$) in [mu2014square, montanari.sun_2018_spectral], we unfold $T$ to an $n\times n^{k-1}$ matrix in the most unbalanced way, and this allows us to directly estimate individual components of the tensor decomposition with sample complexity $O(n^{k/2})$. Note that simply applying the results in [bordenave2020detection] to $\mathrm{unfold}_{\lfloor k/2\rfloor,\lceil k/2\rceil}(T)$ will have sample complexity $O(n^{\lceil k/2\rceil})$, and one cannot directly estimate any individual component in $T$. Therefore, our algorithm has a better computational complexity for tensor completion for all integer values $k$. The long matrix completion result also has a direct application in low-multilinear-rank tensor completion. Detailed statements and discussion on tensor completion are given in Sections LABEL:sec:tensor_completion and LABEL:sec:tucker.
Although there is a strong connection between the detection problem for matrix completion and community detection in the stochastic block model [angelini.caltagirone.ea_2015_spectral, stephan2020non, bordenave2018nonbacktracking, bordenave2020detection], the statistical-to-computational gap in the hypergraph stochastic block model behaves very differently from tensor completion [Pal_2021, stephan2022sparse, gu2023weak]. In hypergraph community detection, efficient spectral methods can achieve weak recovery with the sample complexity down to the information-theoretical threshold up to a constant factor in all cases. However, in tensor completion, there is a conjectured diverging gap between the information-theoretical sample complexity and sample complexity for efficient algorithms [barak.moitra_2016_noisy].
The non-backtracking wedge operator we defined in Section 2.1 is tailored for the situation where the underlying sparse bipartite graph has a diverging aspect ratio. It is similar in spirit to the non-backtracking operator for hypergraphs [angelini.caltagirone.ea_2015_spectral, liu2022statistical, stephan2022sparse], and the nomadic walk operator in [mohanty2020sdp, guruswami2022l_p] for (signed) random bipartite biregular graphs. The main difference compared to [guruswami2022l_p, mohanty2020sdp] is that we work with a non-centered random matrix model without degree homogeneity, and we only count one-sided non-backtracking walks. In addition, a detailed eigenvector analysis is provided while [guruswami2022l_p, mohanty2020sdp] are only about extreme eigenvalues. In terms of tensor estimation with path-counting arguments, another relevant work is the estimation of a spiked tensor model under heavy-tail noise [ding.hopkins.ea_2020_estimating], where the authors use $n\times n$ symmetric matrices to count the number of self-avoiding walks of length $c\log n$. However, those matrices are usually more expensive to compute and require a new parameter (the length of the paths considered) that needs tuning in practice.
2 Setting and main results
2.1 Detailed setting for long matrix completion
We consider a rectangular matrix $M$ of size $n\times m$, with aspect ratio $\alpha=m/n$. We write the singular value decomposition of $M$ as
$M=\sum_{i=1}^{n}\nu_{i}\phi_{i}\psi_{i}^{\top},$
where the $\nu_{i}$ are the singular values of $M$, and the $(\phi_{i})_{i\in[r]}$ (resp. the $(\psi_{i})_{i\in[r]}$) are an orthonormal family of its left (resp. right) singular vectors. We order the singular values of $M$ in decreasing order, that is $\nu_{1}\geq\dots\geq\nu_{n}.$ Our observed matrix is of the form
$A=\frac{\sqrt{mn}}{d}(X\circ M),$ (1)
where $X$ is an i.i.d Bernoulli matrix:
$X_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\ \mathrm{Ber}\left(\frac{d}{\sqrt{mn}}\right).$
We will use the shortcut $p=d/\sqrt{mn}$ as the sampling probability. The scaling of $A$ is chosen so that the expected value of $A$ is exactly $M$. Our goal is to weakly recover the eigendecomposition:
$MM^{\top}=\sum_{i=1}^{n}\nu_{i}^{2}\phi_{i}\phi_{i}^{\top}.$
The non-backtracking wedge matrix
We can view the matrix $A$ as the weighted biadjacency matrix of a bipartite graph $G=(V_{1}\cup V_{2},E)$ with $|V_{1}|=n$, $|V_{2}|=m$. We call $V_{1}$ a left vertex set and $V_{2}$ a right vertex set. A path of length 2 in $G$ denoted by $(e_{1},e_{2},e_{3})$ is called a wedge if $e_{1},e_{3}\in V_{1},e_{2}\in V_{2},$ and $e_{1}\not=e_{3}$. Define the set of oriented wedges $\vec{E}$ as
$\vec{E}=\left\{(e_{1},e_{2},e_{3})\in V_{1}\times V_{2}\times V_{1}:\{e_{1},e_{2}\}\in E,\{e_{2},e_{3}\}\in E,e_{1}\neq e_{3}\right\}.$ (2)
The non-backtracking wedge matrix $B$ associated to $G$ is defined on $\vec{E}\times\vec{E}$ by
$B_{ef}=\mathbf{1}_{e\to f}A_{f_{1}f_{2}}A_{f_{3}f_{2}},\quad\text{where}\quad\mathbf{1}_{e\to f}=\mathbf{1}_{e_{3}=f_{1}}\mathbf{1}_{e_{2}\neq f_{2}}.$ (3)
Equivalently, $\mathbf{1}_{e\to f}$ is one if and only if $(e_{1},e_{2},e_{3},f_{2},f_{3})$ is a non-backtracking walk of length 4 in $G$. See Figure 1 for an example. Note that the non-backtracking wedge matrix $B$ is not symmetric: its index set is the oriented wedge set defined in (2), and $B_{ef}\not=B_{fe}$ in (3).
$e_{1}$$e_{2}$$e_{3}$$f_{2}$$f_{3}$
Figure 1: Example of a non-backtracking path between two wedges $(e_{1},e_{2},e_{3})$ and $(f_{1},f_{2},f_{3})$.
For comparison, we recall the definition of the non-backtracking operator for weighted graphs as follows. Let $A$ be a weighted adjacency matrix of a graph $G=(V,E)$. Define the oriented edge set $\vec{E}$ for $G$ as $\vec{E}=\{(i,j):\{i,j\}\in E\}$. The non-backtracking operator $B$ of associated with $A$ is a non-Hermitian operator of size $|\vec{E}|\times|\vec{E}|$. For any $(u,v),(x,y)\in\vec{E}$, $B$ is defined as
$B_{(u,v),(x,y)}=\begin{cases}A_{xy}&\text{$v=x$, $u\neq y$},\\ 0&\text{otherwise.}\end{cases}$ (4)
In the case of an $n\times m$ bipartite weighted graph defined in (1), the non-backtracking wedge operator only counts non-backtracking walks starting from a vertex in $V_{1}$ while the non-backtracking operator defined in (4) counts non-backtracking walks from both $V_{1}$ and $V_{2}$.
Intuition for the non-backtracking wedge matrix
It has been shown that the non-backtracking operator, as a non-hermitian matrix, outperforms the standard (weighted) adjacency matrix when the sample size is very small in many graph clustering and matrix completion problems [bordenave2018nonbacktracking, bordenave2020detection, stephan2020non, stephan2022sparse]. The key difference is that the spectrum of the non-backtracking operator is more informative and stable when the underlying graph is very sparse with a constant average degree. It is also a general phenomenon that the spectrum of non-hermitian random matrices is less sensitive to a row or column with a large $\ell_{2}$-norm [coste2023sparse, bordenave2022convergence].
In terms of the sample size, the non-backtracking matrix in (4) was used in [bordenave2020detection, stephan2020non] to achieve the optimal scale $O(n)$ for square matrix completion. However, for a low-rank tensor in a matrix form of size $n\times n^{2}$, the work [bordenave2020detection] only achieves weak recovery with sample size $O(n^{2})$, while with the non-backtracking wedge operator, we only need $O(n^{1.5})$. The standard way in [montanari.sun_2018_spectral, cai2021nonconvex] to obtain $\tilde{O}(n^{1.5})$ sample size is to apply a spectral method on the hollowed Gram matrix $Z=AA^{T}-\text{diag}(AA^{T})$. For $i\not=j$, we have
$Z_{ij}=\sum_{\text{wedge }i\to k\to j}A_{ik}A_{kj}.$
Hence, the spectrum of $Z$ depends on wedge walk counts. Our new non-backtracking operator only counts the wedge walks that are non-backtracking, which avoids the localization effect from high-degree vertices [benaychgeorges.bordenave.ea_2019_largest] in the weighted graph corresponding to $Z$. This new operator introduces a tool to “regularize” the spectrum of a heavy-tailed long random matrix without any tuning parameter, which might be useful for other applications.
2.2 Quantifying the convergence
The alignment between the spectral decomposition of $B$ and the SVD of $A$ will be governed by a wealth of different quantities. For ease of understanding, we divide those parameters into two categories, depending on their role in the convergence.
Threshold parameters
These parameters describe which parts of the SVD of $M$ will be reflected in the spectral decomposition of $B$.
  1. (i)
    Variance parameter:
    $\rho=\lVert Q\rVert,\quad\text{where}\quad Q=\sqrt{mn}(M\circ M).$ (5)
  2. (ii)
    Amplitude parameter:
    $L=\sqrt{mn}\max_{x\in[n],y\in[m]}|M_{xy}|.$
  3. (iii)
    Detection thresholds
    $\theta=\max(\theta_{1},\theta_{2}),$
    where the two parameters $\theta_{1}$ and $\theta_{2}$ are defined as
    $\theta_{1}=\sqrt{\frac{\rho}{d}}\quad$ (6)